OpenStudy (anonymous):

a soda pop company determined that there is a 2% chance that a soft drink bottle is under filled if 2 bottles are selected at random find the probability that the 1st one is under filled and the 2nd is not

OpenStudy (ivettef365):

the chance to be under filled = 2/100 the chances not to be = 98/100 so multiply 2/100 x 98/100

OpenStudy (anonymous):

p=0.2 n=2 P1*P2